OpenStudy (anonymous):

Work of 3 Joules done in stretching a spring from its natural length to 18 cm beyond its natural length. What is the force (in Newtons) that holds the spring stretched at the same distance (18 cm)?

OpenStudy (anonymous):

\[W_{spring} = \frac{ 1 }{ 2 }k x^2\]\[F_{spring} = -kx\]

OpenStudy (anonymous):

\[3 Joules = \frac{ 1 }{ 2 }k (0.18)^2\] from this you get k\[F = -k(0.18)\]